OpenStudy (fernandez555):

please identify any errors and explain the corrections that should be made in the following equations: Splott- 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 5x + 4 = 14 5x + 4 + 4 = 14 + 4 5x = 18 5x(divided by)5 = 18(divided by)5 x= 18/5 fizzle: 2x + 4 = –3x + 14 2x + 4 – 2x = –3x + 14 – 2x 4 = 14 – 5x 4 – 14 = 14 – 5x – 14 –12 = –5x –12 + 5 = –5x + 5 –7 = x

OpenStudy (fernandez555):

please point out any errors for each, and correct them- showing first the error, then the correction

OpenStudy (anonymous):

Splott's errors : starting at 4th line.. 5x + 4 - 4 = 14 - 4 5x = 10 x = 2 Fizzle's errors : starting at 5th line.. -10 = -5x -10/-5 = x 2 = x

OpenStudy (anonymous):

can some one pls help me on mine

OpenStudy (fernandez555):

@kelliegirl33 thnk you. and what would the corrections for those errors be?